# Structure-Fair Quantum Circuit Complexity: An Auditable Information-Theoretic Lower Bound

### v2

@@ B.3.1 Analysis of Applicability in Specific Physical Scenarios
Different complexity paradigms have their respective focuses in different physical problems. The unique advantages of RCC lie in its structural fairness and its property as a universal lower bound[40, 38, 39].
@@
    Quantum Error Correction and Logical Computation: When assessing the preparation cost of logical states for different quantum error-correcting codes, a core challenge is to fairly compare different physical encoding schemes [14, 15]. By defining the code space as the reference set, RCC’s calculation result depends only on the complexity of the logical information and is independent of the number of physical qubits [52]. This makes it an ideal, encoding-independent performance benchmark [16]. It can be combined with model-specific resource counts (such as T-count for the Clifford+T model) to provide complementary perspectives [53, 54]. In addition, practical architectures illustrate how logical operations exploit code structure without inflating physical-qubit–level costs [17].
@@
    Hamiltonian Dynamics and Quantum Chaos: When studying the process of state complexification under Hamiltonian evolution, Krylov complexity—through the spreading of an operator in the Krylov basis—provides deep insights into the chaotic nature of the dynamics [9]. RCC can serve as a complementary tool: by selecting a sector of the system’s energy eigenstates as the reference set, one can study the minimum information-theoretic complexity that the final state of the system can achieve under energy constraints, thereby providing a lower bound on the range of states accessible to the dynamical process.
@@
    Quantum Resource Theories: Resource theories, such as those for “mana” or entanglement, often quantify the resource content of a state via relative entropy with respect to a set of free states, $D(\rho\|F)$ [55, 56]. RCC can be viewed as both a special case and a generalization: it concretizes “free” into a reference state $\sigma_{R}$ defined by an arbitrary physical structure and links it directly to computational cost [13]. For stabilizer-based nonclassicality and related resource pricing, representative instances include magic-state measures and operational characterizations [10, 11].
@@ Appendix C Formal Details and Implementation Protocols
This appendix aims to consolidate and formally articulate several key technical details and implementation protocols mentioned in the main text. Its purpose is to provide a deeper level of mathematical support for the rigor of the theory and to offer a precise, operational guide for researchers who wish to apply this framework in experiments or numerical simulations[57, 58].
Base and Shorthand. In this appendix, all logarithms $\log$ are by default base $2$, and $\ln$ denotes the natural logarithm. For the sake of conciseness, let $c=c_{1}/\log\Gamma_{R}>0$, and $D=(D(\rho||\sigma_{R})+\Delta_{\mathrm{spec}}(\rho))/\log\Gamma_{R}-c\cdot\log(1/\varepsilon)$. The main inequality, Eq. (3.1), can be written compactly as $x\geq D-c\log(x)$, and its corresponding equation is $x+c\log(x)=D$.
@@ C.2 Closed-Form Solution and Usable Approximations for the Bootstrap Inequality
The main theorem, Eq. (3.1), leads to a bootstrap-type lower bound $x\geq D-c\log(x)$ (where $x>0$). This section provides its rigorous closed-form minimum solution, its monotonicity and uniqueness, and offers two types of implementable approximations for direct application in numerical or experimental reporting.
@@
$x_{*}=\alpha\cdot W_{0}\left(\frac{1}{\alpha}\cdot\exp\left(\frac{D}{\alpha}\right)\right)$ (C.1)
@@
$x_{*}=D-c\log(D)+o(1)\quad\text{as }D\to\infty$ (C.2)
@@
$x_{*}\geq\begin{cases}\max\{0,D/(1+\alpha)\}&\text{for }0<D\leq 1+\alpha\\ D-\alpha\ln(1+D)&\text{for }D>1+\alpha\end{cases}$ (C.3)
Implementation Recommendation: When performing numerical evaluations, the Lambert-W closed-form solution from Eq. (C.1) should be used preferentially. If a Lambert-W function library is not available, Eq. (C.2) can be used as an initial value for Newton’s method, or the conservative piecewise lower bound given by Eq. (C.3) can be reported directly.
@@ C.3 Robustness to Support Set Leakage: Error Propagation
$D(\rho_{\mathrm{phys}}||\tilde{\sigma}_{R})\geq D(\tilde{\rho}||\sigma_{R})-k\cdot\left(\log\left(\frac{1}{1-p_{\mathrm{leak}}}\right)+\log\left(\frac{1}{\delta}\right)\right)$ (C.4)
Proof Sketch. Let $\rho_{\mathrm{phys}}=q\tilde{\rho}\oplus(1-q)\rho_{\perp}$, where $q=1-p_{\mathrm{leak}}$. Take the smoothed reference state $\tilde{\sigma}_{R}=(1-\delta)\sigma_{R}\oplus\delta\tau_{\perp}$. Consider a binary measure-and-prepare channel $M$ for "whether the state falls within $H_{R}$". According to the data processing inequality and the chain rule for relative entropy, we have $D(\rho_{\mathrm{phys}}||\tilde{\sigma}_{R})\geq D_{\mathrm{Bern}}(q||1-\delta)+qD(\tilde{\rho}||\sigma_{R})$, where $D_{\mathrm{Bern}}$ is the binary classical KL divergence. Using the continuity bound for $D_{\mathrm{Bern}}$, the first term can be lower-bounded by a term of the form $k(\log(1/q)+\log(1/\delta))$, where $k=O(1)$ is a universal constant. From this, Eq. (C.4) follows.
@@ C.4 Standard Relations for One-Shot Entropies
$D_{H}^{\eta}(\rho||\sigma_{R})\leq D_{\max}^{\varepsilon}(\rho||\sigma_{R})+k_{1}\cdot(\log(1/\eta)+\log(1/\varepsilon))$ (C.5)
$D(\rho||\sigma_{R})\geq D_{H}^{\eta}(\rho||\sigma_{R})-k_{2}\cdot\log(1/\eta)$ (C.6)
where all $\log$ are base $2$. These inequalities serve as the bridge connecting Theorem 3.1 and Theorem 3.2, and their proofs rely on standard techniques from one-shot quantum information theory [59].
@@ Appendix D Technical Details and Statistical Protocols for the Operational Principles
This appendix provides the technical details and statistical protocols for the three information extraction principles proposed in Chapter 4 of the main text: quantum hypothesis testing, projective witnesses, and the dephasing entropy bound. Its core objective is to provide testable lower bounds for $D_{H}^{\eta}(\rho||\sigma_{R})$, $D_{\max}(\rho||\sigma_{R})$, or $D(\rho||\sigma_{R})$ with rigorous statistical control at a confidence level of $1-\delta$, so as to facilitate the application of the main theorem to real data. We first unify the notation and prerequisites[57]:
@@ D.1 Quantum Hypothesis Testing Protocol
$D_{H}^{\eta}(\rho||\sigma_{R})=-\log\left(\min_{0\leq T\leq I,\mathrm{Tr}(T\sigma_{R})\leq\eta}\mathrm{Tr}((I-T)\rho)\right)$ (D.1)
@@
$D_{H}^{\eta}(\rho||\sigma_{R})\geq-\log(\beta_{U}(\delta))$ (D.2)
@@ D.2 Projective Witness Protocol
$D_{\max}(\rho||\sigma_{R})\geq\max\{0,\log(p_{L}(\delta)\cdot d_{R}/r)\}$ (D.5)
@@
$N_{\min}\geq\frac{1}{2(p_{0}-p_{\mathrm{star}})^{2}}\cdot\ln(1/\delta)$ (D.6)
@@ D.3 Dephasing Entropy Bound Protocol
$D(\rho||\sigma_{R})\geq D(\Delta_{R}(\rho)||\sigma_{R})=\log(d_{R})-H(P)$ (D.7)
@@
$\varepsilon_{N}(\delta)=\sqrt{\frac{2}{N}\cdot(M\ln 2+\ln(1/\delta))}$ (D.8)
@@ Appendix E Derivation and Physical Meaning of the RECT-$\eta$ Performance Equation
This appendix aims to provide a self-contained and complete derivation for the RECT-$\eta$ performance equation cited in Sec 5.2 of the main text, and to elucidate its nature as a “resource accounting identity”[31][60][61][32, 33].
@@ E.1 Theoretical Foundations: Efficiency Factors and Equalization
The starting point of our theory is two independent, universal inequalities that describe the time cost of a quantum dynamical process; they constitute the axiomatic foundation of the RECT framework.
@@
$\sigma_{\mathrm{avail}}\cdot\Delta_{t}\geq\left(\frac{\pi\hbar}{2}\right)\cdot C_{\mathrm{opt}}$ (E.1)
@@
$S_{E}\leq\left(\frac{\gamma J}{\hbar}\right)\cdot\Delta_{t}$ (E.2)
@@
$\eta_{\mathrm{QSL}}=\frac{(\pi\hbar/2)\cdot C_{\mathrm{opt}}}{\sigma_{\mathrm{avail}}\cdot\Delta_{t}}$ (E.3)
$\eta_{\mathrm{LR}}=\frac{S_{E}}{(\gamma J/\hbar)\cdot\Delta_{t}}$ (E.4)
According to Eq. (E.1) and Eq. (E.2), the physical value ranges for these two efficiency factors are $\eta_{\mathrm{QSL}}\in(0,1]$ and $\eta_{\mathrm{LR}}\in[0,1]$, respectively. These two definitions precisely convert the aforementioned inequalities into equalities:
$\sigma_{\mathrm{avail}}\cdot\Delta_{t}=\frac{(\pi\hbar/2)\cdot C_{\mathrm{opt}}}{\eta_{\mathrm{QSL}}}$ (E.5)
$S_{E}=\eta_{\mathrm{LR}}\cdot\left(\frac{\gamma J}{\hbar}\right)\cdot\Delta_{t}$ (E.6)
@@ E.2 Core Derivation
Eq. (E.5) and Eq. (E.6) provide two exact expressions for the same time interval $\Delta_{t}$ of the same physical process, which allows us to eliminate $\Delta_{t}$ through algebraic methods. Solving for $\Delta_{t}$ from both equations separately (assuming $S_{E}>0$):
$\Delta_{t}=\frac{\pi\hbar C_{\mathrm{opt}}}{2\sigma_{\mathrm{avail}}\eta_{\mathrm{QSL}}}$ (E.7)
$\Delta_{t}=\frac{\hbar S_{E}}{\eta_{\mathrm{LR}}\gamma J}$ (E.8)
@@
$\frac{\pi\hbar C_{\mathrm{opt}}}{2\sigma_{\mathrm{avail}}\eta_{\mathrm{QSL}}}=\frac{\hbar S_{E}}{\eta_{\mathrm{LR}}\gamma J}$ (E.9)
@@
$\frac{\pi C_{\mathrm{opt}}}{2\sigma_{\mathrm{avail}}\eta_{\mathrm{QSL}}}=\frac{S_{E}}{\eta_{\mathrm{LR}}\gamma J}$ (E.10)
@@
$\sigma_{\mathrm{avail}}\cdot S_{E}=\left(\frac{\eta_{\mathrm{LR}}}{\eta_{\mathrm{QSL}}}\right)\cdot\left(\frac{\pi\gamma J}{2}\right)\cdot C_{\mathrm{opt}}$ (E.11)
@@ E.3 Physical Essence: The Identity and Testable Content
Proposition E.1 (The Identity Nature of RECT-$\eta$): Eq. (E.11), by its construction, is an algebraic identity.
Proof. Substitute the definitions of the efficiency factors, Eq. (E.3) and Eq. (E.4), into the right-hand side of Eq. (E.11):
@@
$\mathrm{RHS}$ $=\frac{S_{E}\cdot\sigma_{\mathrm{avail}}\cdot\Delta_{t}}{\bigl(\tfrac{\gamma J}{\hbar}\bigr)\Delta_{t}\cdot\bigl(\tfrac{\pi\hbar}{2}\bigr)C_{\mathrm{opt}}}\cdot\left(\frac{\pi\gamma J}{2}\right)\cdot C_{\mathrm{opt}}$ (E.12)
$=\frac{S_{E}\cdot\sigma_{\mathrm{avail}}}{\tfrac{\gamma J\pi\hbar^{2}}{2\hbar}}\cdot\left(\frac{\pi\gamma J}{2}\right)\qquad(\text{note: }\hbar^{2}/\hbar=\hbar)$ (E.13)
$=S_{E}\cdot\sigma_{\mathrm{avail}}.$ (E.14)
@@
The value of the RECT-$\eta$ equation lies not in testing the equality itself, but in the fact that it provides a unified “resource accounting framework”. Within this framework, two seemingly unrelated physical laws are connected, allowing us to define a powerful set of diagnostic tools for dynamics (such as the efficiency ratio $\eta_{\mathrm{LR}}/\eta_{\mathrm{QSL}}$) and performance benchmarks (such as the performance frontier) to conduct in-depth analysis of real quantum processes.For the detailed derivation and discussion of this performance equation, please refer to the original paper [31].

### v3

@@ B.3.1 Analysis of Applicability in Specific Physical Scenarios
Different complexity paradigms have their respective focuses in different physical problems. The unique advantages of RCC lie in its structural fairness and its property as a universal lower bound[Howard2014, NielsenGeometry06, parker2019universal].
@@
    Quantum Error Correction and Logical Computation: When assessing the preparation cost of logical states for different quantum error-correcting codes, a core challenge is to fairly compare different physical encoding schemes [Gottesman1997, Fowler2012]. By defining the code space as the reference set, RCC’s calculation result depends only on the complexity of the logical information and is independent of the number of physical qubits [BravyiHaah2012]. This makes it an ideal, encoding-independent performance benchmark [BravyiKitaev2005]. It can be combined with model-specific resource counts (such as T-count for the Clifford+T model) to provide complementary perspectives [Amy2013, Selinger2013]. In addition, practical architectures illustrate how logical operations exploit code structure without inflating physical-qubit–level costs [GidneyFowler2019].
@@
    Hamiltonian Dynamics and Quantum Chaos: When studying the process of state complexification under Hamiltonian evolution, Krylov complexity—through the spreading of an operator in the Krylov basis—provides deep insights into the chaotic nature of the dynamics [parker2019universal]. RCC can serve as a complementary tool: by selecting a sector of the system’s energy eigenstates as the reference set, one can study the minimum information-theoretic complexity that the final state of the system can achieve under energy constraints, thereby providing a lower bound on the range of states accessible to the dynamical process.
@@
    Quantum Resource Theories: Resource theories, such as those for “mana” or entanglement, often quantify the resource content of a state via relative entropy with respect to a set of free states, $D(\rho\|F)$ [Vedral1997, BrandaoPlenio2010]. RCC can be viewed as both a special case and a generalization: it concretizes “free” into a reference state $\sigma_{R}$ defined by an arbitrary physical structure and links it directly to computational cost [BrandaoGour2015]. For stabilizer-based nonclassicality and related resource pricing, representative instances include magic-state measures and operational characterizations [Veitch2014, Howard2014].
@@ Appendix C Formal Details and Implementation Protocols
This appendix aims to consolidate and formally articulate several key technical details and implementation protocols mentioned in the main text. Its purpose is to provide a deeper level of mathematical support for the rigor of the theory and to offer a precise, operational guide for researchers who wish to apply this framework in experiments or numerical simulations[Tomamichel2016Finite, Watrous2018].
Base and Shorthand. In this appendix, all logarithms $\log$ are by default base $2$, and $\ln$ denotes the natural logarithm. For the sake of conciseness, let $c=c_{1}/\log\Gamma_{R}>0$, and $D=(D(\rho||\sigma_{R})+\Delta_{\mathrm{spec}}(\rho))/\log\Gamma_{R}-c\cdot\log(1/\varepsilon)$. The main inequality, Eq. (6), can be written compactly as $x\geq D-c\log(x)$, and its corresponding equation is $x+c\log(x)=D$.
@@ C.2 Closed-Form Solution and Usable Approximations for the Bootstrap Inequality
The main theorem, Eq. (6), leads to a bootstrap-type lower bound $x\geq D-c\log(x)$ (where $x>0$). This section provides its rigorous closed-form minimum solution, its monotonicity and uniqueness, and offers two types of implementable approximations for direct application in numerical or experimental reporting.
@@
$x_{*}=\alpha\cdot W_{0}\left(\frac{1}{\alpha}\cdot\exp\left(\frac{D}{\alpha}\right)\right)$ (56)
@@
$x_{*}=D-c\log(D)+o(1)\quad\text{as }D\to\infty$ (57)
@@
$x_{*}\geq\begin{cases}\max\{0,D/(1+\alpha)\}&\text{for }0<D\leq 1+\alpha\\ D-\alpha\ln(1+D)&\text{for }D>1+\alpha\end{cases}$ (58)
Implementation Recommendation: When performing numerical evaluations, the Lambert-W closed-form solution from Eq. (56) should be used preferentially. If a Lambert-W function library is not available, Eq. (57) can be used as an initial value for Newton’s method, or the conservative piecewise lower bound given by Eq. (58) can be reported directly.
@@ C.3 Robustness to Support Set Leakage: Error Propagation
$D(\rho_{\mathrm{phys}}||\tilde{\sigma}_{R})\geq D(\tilde{\rho}||\sigma_{R})-k\cdot\left(\log\left(\frac{1}{1-p_{\mathrm{leak}}}\right)+\log\left(\frac{1}{\delta}\right)\right)$ (59)
Proof Sketch. Let $\rho_{\mathrm{phys}}=q\tilde{\rho}\oplus(1-q)\rho_{\perp}$, where $q=1-p_{\mathrm{leak}}$. Take the smoothed reference state $\tilde{\sigma}_{R}=(1-\delta)\sigma_{R}\oplus\delta\tau_{\perp}$. Consider a binary measure-and-prepare channel $M$ for "whether the state falls within $H_{R}$". According to the data processing inequality and the chain rule for relative entropy, we have $D(\rho_{\mathrm{phys}}||\tilde{\sigma}_{R})\geq D_{\mathrm{Bern}}(q||1-\delta)+qD(\tilde{\rho}||\sigma_{R})$, where $D_{\mathrm{Bern}}$ is the binary classical KL divergence. Using the continuity bound for $D_{\mathrm{Bern}}$, the first term can be lower-bounded by a term of the form $k(\log(1/q)+\log(1/\delta))$, where $k=O(1)$ is a universal constant. From this, Eq. (59) follows.
@@ C.4 Standard Relations for One-Shot Entropies
$D_{H}^{\eta}(\rho||\sigma_{R})\leq D_{\max}^{\varepsilon}(\rho||\sigma_{R})+k_{1}\cdot(\log(1/\eta)+\log(1/\varepsilon))$ (60)
$D(\rho||\sigma_{R})\geq D_{H}^{\eta}(\rho||\sigma_{R})-k_{2}\cdot\log(1/\eta)$ (61)
where all $\log$ are base $2$. These inequalities serve as the bridge connecting Theorem 3.1 and Theorem 3.2, and their proofs rely on standard techniques from one-shot quantum information theory [berta2016fidelity].
@@ Appendix D Technical Details and Statistical Protocols for the Operational Principles
This appendix provides the technical details and statistical protocols for the three information extraction principles proposed Sec. 4 of the main text: quantum hypothesis testing, projective witnesses, and the dephasing entropy bound. Its core objective is to provide testable lower bounds for $D_{H}^{\eta}(\rho||\sigma_{R})$, $D_{\max}(\rho||\sigma_{R})$, or $D(\rho||\sigma_{R})$ with rigorous statistical control at a confidence level of $1-\delta$, so as to facilitate the application of the main theorem to real data. We first unify the notation and prerequisites[Tomamichel2016Finite]:
@@ D.1 Quantum Hypothesis Testing Protocol
$D_{H}^{\eta}(\rho||\sigma_{R})=-\log\left(\min_{0\leq T\leq I,\mathrm{Tr}(T\sigma_{R})\leq\eta}\mathrm{Tr}((I-T)\rho)\right)$ (63)
@@
$D_{H}^{\eta}(\rho||\sigma_{R})\geq-\log(\beta_{U}(\delta))$ (64)
@@ D.2 Projective Witness Protocol
$D_{\max}(\rho||\sigma_{R})\geq\max\{0,\log(p_{L}(\delta)\cdot d_{R}/r)\}$ (67)
@@
$N_{\min}\geq\frac{1}{2(p_{0}-p_{\mathrm{star}})^{2}}\cdot\ln(1/\delta)$ (68)
@@ D.3 Dephasing Entropy Bound Protocol
$D(\rho||\sigma_{R})\geq D(\Delta_{R}(\rho)||\sigma_{R})=\log(d_{R})-H(P)$ (69)
@@
$\varepsilon_{N}(\delta)=\sqrt{\frac{2}{N}\cdot(M\ln 2+\ln(1/\delta))}$ (70)
@@ Appendix E Derivation and Physical Meaning of the RECT-$\eta$ Performance Equation
This appendix aims to provide a self-contained and complete derivation for the RECT-$\eta$ performance equation cited in Sec V.2 of the main text, and to elucidate its nature as a “resource accounting identity”[liu2025cost][Pandey2023SpeedLimits].
@@ E.1 Theoretical Foundations: Efficiency Factors and Equalization
The starting point of our theory is two independent, universal inequalities that describe the time cost of a quantum dynamical process; they constitute the axiomatic foundation of the RECT framework.[Pandit2022boundsrecurrence][MandelstamTamm1945, LiebRobinson1972]
@@
$\sigma_{\mathrm{avail}}\cdot\Delta_{t}\geq\left(\frac{\pi\hbar}{2}\right)\cdot C_{\mathrm{opt}}$ (73)
@@
$S_{E}\leq\left(\frac{\gamma J}{\hbar}\right)\cdot\Delta_{t}$ (74)
@@
$\eta_{\mathrm{QSL}}=\frac{(\pi\hbar/2)\cdot C_{\mathrm{opt}}}{\sigma_{\mathrm{avail}}\cdot\Delta_{t}}$ (75)
$\eta_{\mathrm{LR}}=\frac{S_{E}}{(\gamma J/\hbar)\cdot\Delta_{t}}$ (76)
According to Eq. (73) and Eq. (74), the physical value ranges for these two efficiency factors are $\eta_{\mathrm{QSL}}\in(0,1]$ and $\eta_{\mathrm{LR}}\in[0,1]$, respectively. These two definitions precisely convert the aforementioned inequalities into equalities:
$\sigma_{\mathrm{avail}}\cdot\Delta_{t}=\frac{(\pi\hbar/2)\cdot C_{\mathrm{opt}}}{\eta_{\mathrm{QSL}}}$ (77)
$S_{E}=\eta_{\mathrm{LR}}\cdot\left(\frac{\gamma J}{\hbar}\right)\cdot\Delta_{t}$ (78)
@@ E.2 Core Derivation
Eq. (77) and Eq. (78) provide two exact expressions for the same time interval $\Delta_{t}$ of the same physical process, which allows us to eliminate $\Delta_{t}$ through algebraic methods. Solving for $\Delta_{t}$ from both equations separately (assuming $S_{E}>0$):
$\Delta_{t}=\frac{\pi\hbar C_{\mathrm{opt}}}{2\sigma_{\mathrm{avail}}\eta_{\mathrm{QSL}}}$ (79)
$\Delta_{t}=\frac{\hbar S_{E}}{\eta_{\mathrm{LR}}\gamma J}$ (80)
@@
$\frac{\pi\hbar C_{\mathrm{opt}}}{2\sigma_{\mathrm{avail}}\eta_{\mathrm{QSL}}}=\frac{\hbar S_{E}}{\eta_{\mathrm{LR}}\gamma J}$ (81)
@@
$\frac{\pi C_{\mathrm{opt}}}{2\sigma_{\mathrm{avail}}\eta_{\mathrm{QSL}}}=\frac{S_{E}}{\eta_{\mathrm{LR}}\gamma J}$ (82)
@@
$\sigma_{\mathrm{avail}}\cdot S_{E}=\left(\frac{\eta_{\mathrm{LR}}}{\eta_{\mathrm{QSL}}}\right)\cdot\left(\frac{\pi\gamma J}{2}\right)\cdot C_{\mathrm{opt}}$ (83)
@@ E.3 Physical Essence: The Identity and Testable Content
Proposition E.1 (The Identity Nature of RECT-$\eta$): Eq. (83), by its construction, is an algebraic identity.
Proof. Substitute the definitions of the efficiency factors, Eq. (75) and Eq. (76), into the right-hand side of Eq. (83):
@@
$\mathrm{RHS}$ $=\frac{S_{E}\,\sigma_{\mathrm{avail}}\,\Delta_{t}}{\bigl(\tfrac{\gamma J}{\hbar}\bigr)\Delta_{t}\,\bigl(\tfrac{\pi\hbar}{2}\bigr)C_{\mathrm{opt}}}\left(\frac{\pi\gamma J}{2}\right)C_{\mathrm{opt}}$ (84)
$=\frac{S_{E}\,\sigma_{\mathrm{avail}}}{\bigl(\tfrac{\gamma J}{\hbar}\bigr)\bigl(\tfrac{\pi\hbar}{2}\bigr)}\left(\frac{\pi\gamma J}{2}\right)\qquad\text{(note: }\hbar^{2}/\hbar=\hbar\text{)}$
$=S_{E}\,\sigma_{\mathrm{avail}}.$
@@
The value of the RECT-$\eta$ equation lies not in testing the equality itself, but in the fact that it provides a unified “resource accounting framework”. Within this framework, two seemingly unrelated physical laws are connected, allowing us to define a powerful set of diagnostic tools for dynamics (such as the efficiency ratio $\eta_{\mathrm{LR}}/\eta_{\mathrm{QSL}}$) and performance benchmarks (such as the performance frontier) to conduct in-depth analysis of real quantum processes.For the detailed derivation and discussion of this performance equation, please refer to the original paper [liu2025cost].